Show that the maximum value of a 2 b 2 c 2 on a sphere of radius r centered at the origin of a Cartesian abc-coordinate system in (r 2/3)3 .


1
Expert's answer
2021-11-03T18:04:51-0400

Let "f(a,b,c)=a^2b^2c^2"

Where "a^2+b^2+c^2=r^2"

"\implies g(a,b,c)=a^2+b^2+c^2"

Use Lagrange multipliers to find the maximum value of "a^2b^2c^2"

"\nabla f(a,b,c)=\lambda\nabla g"

"\implies<2ab^2c^2,2a^2bc^2,2a^2b^2c>=\lambda<2a,2b,2c>"

"\implies ab^2c^2=a\lambda" , "a^2bc^2=b\lambda" , "a^2b^2c=c\lambda"

"\implies \lambda=b^2c^2" , "\lambda=a^2c^2" , "\lambda=a^2b^2"

"\implies b^2c^2=a^2c^2=a^2b^2\implies a=b=c"

Substitute a=b=c in "a^2+b^2+c^2=r^2"

"a=b=c=\frac{r}{\sqrt3}"

"(a^2b^2c^2)_{max}=(\frac{r^2}{3})^3"